OpenStudy (anonymous):

•How can you represent the terms of a sequence? •How can you define an arithmetic or geometric sequence? •How can you model the sum of an arithmetic or geometric series?

OpenStudy (loser66):

take a look at this , everything is there, http://www.purplemath.com/modules/series.htm short answer for this arithmetic : the next term = previous term +/- a constant. for example : 1,3,5,.... ( 3=1+2, 5 =3+2 ....) geometric: next term = previous term * (time) a constant, for example 2,4,8,16..... 4 =2*2 8=4*2 16=8*2 got it?